OpenStudy (anonymous):

Find the solutions of the equation. 1/2x^2-x+5=0

OpenStudy (anonymous):

Use the general quadratic solution formula.

OpenStudy (anonymous):

ax² + bx + c = 0 ?

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two rots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

How do I find a,b&c?

hartnn (hartnn):

 now can u find a,b,c ?